Multiple Choice Question 113 A company is considering purchasing factory equipment which costs $500000 and is estimated to have no salvage value at the end of its 8-year useful life. If the equipment is purchased, annual revenues are expected to be $229000 and annual operating expenses exclusive of depreciation expense are expected to be $90000. The straight-line method of depreciation would be used. If the equipment is purchased, the annual rate of return expected on this project is
Answer #1

Calculation of Annual Rate of Return:

Annual Rate of Return = Net income/Cost of investment

= 256,500/500,000

= 51.3%

Calculation of Net income:

Net income = Revenue-Depreciation expense-other expenses

= 229,000-62,500+90,000

= 256,500

Depreciation expense = Cost of asset-salvage value/Useful life

= 500,000-0/8

= 62,500